OpenStudy (anonymous):

a circle is shown on the coordinate plane below. what is the equation of this circle

OpenStudy (anonymous):

What is the equation of this circle? (x + 1)2 + y2 = 2 (x – 1)2 + y2 = 4 (x + 1)2 + y2 = 4 (x – 1)2 + y2 = 2

OpenStudy (anonymous):

equation of circle is given by formula: \[\left( x-h \right)^{2} + \left( y-h \right)^{2} = r ^{2}\] where (h,k) is the center and r is the radius. in your case the center is (-1,0) and r is \[\left| 1--3 \right|\] substitute in \[\left( x-h \right)^{2} + \left( y-h \right)^{2} = r ^{2}\] it becomes \[\left( x--1 \right)^{2} + \left( y-0 \right)^{2} = 2^{2}\] now simplify

OpenStudy (anonymous):

note the \[\left| 1--3 \right|\] is the diameter not radius